OpenStudy (anonymous):

V^2=2gh use calculus, if g increase by 5% and h decrease by 10% find the percentage change in V

zepdrix (zepdrix):

@ganeshie8 hmm

OpenStudy (anonymous):

??

OpenStudy (anonymous):

calculus?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then i must be reading this wrong

OpenStudy (anonymous):

take log and differentiate dh is delta H the you can use percentages

OpenStudy (anonymous):

then*

OpenStudy (anonymous):

i get \[V^2=2(1.05)g(1.1)h\]

OpenStudy (anonymous):

can you do it out

OpenStudy (anonymous):

because i really dont know how to write it out

OpenStudy (anonymous):

if g increase by 5%\[g\to 1.05g\]

OpenStudy (anonymous):

yep but how to do it in calculus

OpenStudy (anonymous):

i think i do not understand this question, so i will be quiet

OpenStudy (anonymous):

oooh maybe it is this \[2VV'=2(hh'+gg')\] then plug in \(h'=.05,g'=1.1\) does that seem reasonable?

OpenStudy (welshfella):

g is decreasing so it think that would be g'= 0.90

OpenStudy (anonymous):

h0.9 and g 1.05?

OpenStudy (welshfella):

sorry h' = 0.90

OpenStudy (anonymous):

oh right i read it wrong sorry

OpenStudy (anonymous):

hmm not sure about that though, since \(h'\) is the rate of change in my original method, yes, but here i guess \(h'=-.1\)

OpenStudy (anonymous):

2 log V = log g + log h differentiating \[ 2\Delta V / V = \Delta g /g + \Delta h /h\]

OpenStudy (anonymous):

del g /g = 0.05 and del h / h = 0.1